OpenStudy (anonymous):

Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

Nnesha (nnesha):

you should set binomial equal to 0 and then solve that for giving variable then replace polynomial variable by that number if you get 0 then binomial is a factor of polynomial

Nnesha (nnesha):

for example if binomial is ( x+1) set it equal to 0 \[\huge\rm x+1 = 0\] solve for x

Nnesha (nnesha):

polynomial is \[\huge\rm x^2 +4x +3\] replace x b its value if you get 0 then (x+1) is a factor of (x^2 +4x+3)

OpenStudy (anonymous):

(x^2+2x+1)(x-4) how would you solve this ?

OpenStudy (anonymous):

I'm confused on how to solve these @Nnesha

Nnesha (nnesha):

do you have to divide by x-4 ?

Nnesha (nnesha):

divide polynomial by x-4 ??***

OpenStudy (anonymous):

I'm not sure.. we can skip that could you just help me with the example you gave me? X would equal -1 for the binomial

Nnesha (nnesha):

yes right now replace x by -1 in x^2 +4x +3 <--- in this polynomial if you get 0 then x +1 is a factor of x^2 +4x +3

OpenStudy (anonymous):

yes I did get 0

Nnesha (nnesha):

yes then x+1 is a factor of x^2 +4x + 3

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @ayyyyyyyyy25 (x^2+2x+1)(x-4) how would you solve this ? \(\color{blue}{\text{End of Quote}}\) do you have to divide or find if (x-4) is a factor of a polynomial ?

OpenStudy (anonymous):

for (x^2+2x+1)(x-4) it says to just multiply it out @Nnesha

OpenStudy (anonymous):

I think you divide im not sure

Nnesha (nnesha):

\[\huge\rm (x^2 +2x +1 )( x-4) = 0\] or \[ \huge\rm \frac{ (x^2 +2x +1 ) }{ (x-4) }\]